OpenStudy (kewlgeek555):

What steps do I take to transform x^2 + 4x + 3 = 0 into the form (x - p)^2 = q?

mathslover (mathslover):

Use the method of Square Completion here.

OpenStudy (johnweldon1993):

Complete the square

mathslover (mathslover):

\(x^2 + 2(2)x + 4-4 + 3 =0 \) \((x^2 + 2(2)x + 4 ) + 3 - 4= 0\) What I did is that, I assumed 4x as the term equivalent to : 2xp and so I made it like : 2(2)(x) So, we have , x =x and p = 2 We know the formula : \((x-p)^2 = x^2 - 2xp + p^2 \) We have, x^2 and 2xp but we require p^2 So, we add and subtract 2^2 in the LHS That is : (the first step in the first Equation)

mathslover (mathslover):

Can you do it from here?

OpenStudy (johnweldon1993):

Step 1) Subtract 3 from both sides. Step 2) Take the coefficient of the 'x' ..divide it by 2....square the result...and add to both sides of the equation \[\large x^2+4x=−3\] 4/2 = 2 ...2^2 = 4...so we have \[\large x^2+4x+4=−3+4\] Factor the left hand side and simplify the right hand side ...what do you get?

OpenStudy (kewlgeek555):

I am a bit confused @mathslover . (>.<) Wasn't the Square Completion method where first you move the loose number to the other side of the equation, take half of the x-term and square it and then add the square to both sides of the equation...actually, I think the method I learned was the one that @johnweldon1993 seems to be showing...it is refreshing my memory.

mathslover (mathslover):

It depends on your understanding actually, if you think John's method is what you have studied, then follow it... Both methods are correct

OpenStudy (johnweldon1993):

^Indeed ...I have used @mathslover method before as well ...both are correct :)

OpenStudy (kewlgeek555):

Oh, okay, thanks! :D So, let me factor the left hand side... (x + 2)(x + 2) And to simplify the right hand ride would be...1. So (x + 2)(x + 2) = 1 or...\[(x+2)^2=1\]

OpenStudy (johnweldon1993):

Correct @kewlgeek555

OpenStudy (kewlgeek555):

Yay. \o/ Thank you both. ;) I should read the actual lesson and not just the review. (/)v(\)

OpenStudy (johnweldon1993):

Yeah that might help lol...and no problem!

OpenStudy (kewlgeek555):

Also, I probably have a few more, but the most important is one that I want to check if I did it right, I can tag you guys in a new question. cx